REPLACEMENT ANALYSIS

The Bigbee Bottling Company is contemplating the replacement of one of its bottling machines with a newer and more efficient one. The old machine has a book value of $650,000 and a remaining useful life of 5 years. The firm does not expect to realize any return from scrapping the old machine in 5 years, but it can sell it now to another firm in the industry for $280,000. The old machine is being depreciated by $130,000 per year, using the straight-line method. The new machine has a purchase price of $1,150,000, an estimated useful life and MACRS class life of 5 years, and an estimated salvage value of $135,000. The applicable depreciation rates are 20%, 32%, 19%, 12%, 11%, and 6%. It is expected to economize on electric power usage, labor, and repair costs, as well as to reduce the number of defective bottles. In total, an annual savings of $250,000 will be realized if the new machine is installed. The company's marginal tax rate is 35%, and it has a 12% WACC.

a. What initial cash outlay is required for the new machine? Round your answer to the nearest dollar. Negative amount should be indicated by a minus sign. $

b. Calculate the annual depreciation allowances for both machines and compute the change in the annual depreciation expense if the replacement is made. Round your answers to the nearest dollar.

Year 1,2,3,4,5 Depreciation Allowance, New Depreciation Allowance, Old Change in Depreciation. What are the incremental net cash flows in Years 1 through 5? Round your answers to the nearest dollar.

c. Year 1 Year 2 Year 3 Year 4 Year 5

d. Should the firm purchase the new machine?
Answer #1
a] Cost of the new machine $    11,50,000
Current sale value of the old machine $         2,80,000
Book value of the old machine $         6,50,000
Loss on sale of old machine = 650000-280000 = $         3,70,000
Tax shield on loss = 370000*35% = $         1,29,500
After tax sale proceeds of old machine = 280000+129500 = $      4,09,500
Initial cash outlay = 1150000-409500 = $      7,40,500
b] 1 2 3 4 5
New depreciation allowance $      2,30,000 $      3,68,000 $           2,18,500 $        1,38,000 $    1,26,500 $        69,000 $ 11,50,000
Old depreciation allowance $      1,30,000 $      1,30,000 $           1,30,000 $        1,30,000 $    1,30,000
Change in depreciation $      1,00,000 $      2,38,000 $               88,500 $              8,000 $        -3,500
c] Annual savings $      2,50,000 $      2,50,000 $           2,50,000 $        2,50,000 $    2,50,000
-Change in depreciation $      1,00,000 $      2,38,000 $               88,500 $              8,000 $        -3,500
=Incremental NOI $      1,50,000 $          12,000 $           1,61,500 $        2,42,000 $    2,53,500
-Tax at 35% $          52,500 $            4,200 $               56,525 $           84,700 $        88,725
=Incremental NOPAT $          97,500 $            7,800 $           1,04,975 $        1,57,300 $    1,64,775
+Change in depreciation $      1,00,000 $      2,38,000 $               88,500 $              8,000 $        -3,500
=Incremental OCF $      1,97,500 $      2,45,800 $           1,93,475 $        1,65,300 $    1,61,275
-Initial outlay $         7,40,500
+After tax salvage value [See calculation below] $    1,11,900
Incremental net cash flows $        -7,40,500 $      1,97,500 $      2,45,800 $           1,93,475 $        1,65,300 $    2,73,175
PVIF at 12% [PVIF = 1/1.12^t] 1 0.89286 0.79719 0.71178 0.63552 0.56743
PV at 12% $        -7,40,500 $      1,76,339 $      1,95,950 $           1,37,712 $        1,05,051 $    1,55,007
NPV $             29,559
d] As the NPV of the replacement is positive, the firm should purchase the new machine.
CALCULATION OF AFTER TAX SALVAGE VALUE OF NEW MACHINE AT EOY 5:
Salvage value $       $ 1,35,000
-Book value $        $     69,000
=Gain on sale $        $     66,000
Tax on gain = 66000*5% = $      $ 23,100
After tax salvage value = 135000-23100 = $      $ 1,11,900